In a certain country, license plates have three letters followed by two digits. If the first digit cannot be zero, how many different license plates are possible if: repetition of letters are NOT allowed?


1
Expert's answer
2021-03-03T05:55:36-0500

Possible 3 digits combinations: "D=9*10*10=900."

Possible 2 letters combinations: "L=26*25=650."

Number of different license plates: "N=L*D=650*900=585000."